OpenStudy (anonymous):

what is the y-intercept of the lne that has a slope of -1/2 and passes through the point (2,3)?

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &(2\quad ,&3) \end{array} \\\quad \\ slope = m= -\cfrac{1}{2} \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values and solve for "y"}\)

OpenStudy (anonymous):

ok, so y-3=-1/2(x-2)?